OpenStudy (azionne):

12(x - 6) = 2(x - 9) medal

pooja195 (pooja195):

Start by distributing  What equation do you get?

OpenStudy (azionne):

12x-72=2x-18

OpenStudy (texaschic101):

now you want to get all the x terms on one side and everything else on the other side. so first, subtract 2x from both sides and what do you get ?

pooja195 (pooja195):

Good now subtract 2x from both sides \[\huge~\rm~12x-2x-72=2x-2x-18\]

OpenStudy (texaschic101):

now simplify both sides and what do you get ?

pooja195 (pooja195):

@texaschic101 they never said what would be left once we subtract 2x from both sides.

OpenStudy (texaschic101):

simplify both sides and that is what you have left..before the next step

pooja195 (pooja195):

\[\huge~\rm~10x-72=-18\] Still need to add 72 to both sides \[\huge~\rm~10x-72+72=-18+72\]

OpenStudy (texaschic101):

yep...thats the next step

OpenStudy (texaschic101):

now simplify both sides again....then you just have 1 more step to go

OpenStudy (azionne):

i dont get it

pooja195 (pooja195):

What dont you get?

OpenStudy (azionne):

nvm